3.  A number is selected randomly from a container containing all the integers from 10 to 50. Find:

  1. P(even | greater than 40)
  2. P(greater than 40 | even)
  3. P(prime | between 20 and 40)
1
Expert's answer
2022-03-07T05:21:03-0500

1.

Numbers greater than 40:"41,42,43,44,45,46,47,48,49,50"


"P(even|greater\ than\ 40)=\dfrac{5}{10}=\dfrac{1}{2}"

2. There are 25 even numbers.

Numbers greater than 40:"41,42,43,44,45,46,47,48,49,50"


"P(even|greater\ than\ 40)=\dfrac{5}{10}=\dfrac{1}{2}"


Even numbers greater than 40: "42,44,46,48,50"


"P(greater\ than\ 40|even)=\dfrac{5}{25}=\dfrac{1}{5}"

3. There are 19 numbers between 20 and 40.

Prime numbers between 20 and 40: "23,29,31,37."


"P(prime|between\ 20\ and\ 40)=\dfrac{4}{19}"